# Stylos: Multi-View 3D Stylization with Single-Forward Gaussian Splatting

### v1

@@ A.3 Additional Visual Results
Figure 7-16 presents additional qualitative results of Stylos on diverse style–content pairs. Across different styles and object categories, Stylos produces visually pleasing stylizations that respect the input content geometry while transferring the target artistic appearance. These examples further demonstrate the versatility and robustness of our approach beyond the main results shown in the paper.
@@
Figure 7: Additional visual results of Stylos on diverse style–content pairs on Tanks and Temples Dataset (Knapitsch et al. (2017a)). Stylos consistently produces visually pleasing stylizations that preserve content geometry while transferring the target style.

### v2

@@ A.6 Additional Visual Results
Figure 14-15 exhibit all the 50 test style images for quantitative comparisons. Figure 16-25 presents additional qualitative results of Stylos on diverse style–content pairs. Across different styles and object categories, Stylos produces visually pleasing stylizations that respect the input content geometry while transferring the target artistic appearance. These examples further demonstrate the versatility and robustness of our approach beyond the main results in the paper.
@@
Figure 16: Additional visual results of Stylos on diverse style–content pairs on Tanks and Temples Dataset (Knapitsch et al., 2017). Stylos consistently produces visually pleasing stylizations that preserve content geometry while transferring the target style.